# On the Moduli Space of Discrete Morse Functions

### v1

@@ 4. Relationship to smooth Morse theory
[…]
Before we prove this proposition, we need two lemmas.
Lemma 4.5.
Let $f_{t}\colon I\rightarrow\mathcal{M}_{smo}(M)$ be a path in the space of Morse functions [^7] Sometimes called a 1-parameter family of Morse functions.. Then
  1. (1)
    the critical points of $f_{t}$, i.e. the zeros of $Df_{t}$ depend continuously on $t$, and
  2. (2)
    for any $p\in M$ the eigenvalues of $Hf_{t}(p)$ depend continuously on $t$.
Proof.
  1. (1)
    Since $f_{t}$ depends continuously on $t$ and since for any local coordinates the entries of the Jacobian $Df_{t}(p)$ at any point $p\in M$ are the first partial derivatives of $f_{t}$ at $p$, the entries of $Df_{t}$ depend continuously on $t$ because $t$ is just a parameter for the partial derivatives. By definition, $p$ is a critical point of $f$ if and only if $\operatorname{det}Df_{t}(p)\neq 0$. The statement follows because the determinant is continuous in the entries of the matrix.
  2. (2)
    Similar to (1) the entries of the Hessian $Hf_{t}(p)$ at any point $p$ depend continuously on $t$. Since $Hf_{t}(p)$ is symmetric, all its eigenvalues are real. Since the eigenvalues of $Hf_{t}(p)$ are zeros of the characteristic polynomial $\chi_{Hf_{t}(p)}$ and the coefficients of $\chi_{Hf_{t}(p)}$ depend continuously on the entries of $Hf_{t}(p)$, the statement follows directly by [Tyr97, Theorem 3.9.1].
Lemma 4.6.
Along any path $f_{t}\colon I\rightarrow\mathcal{MS}_{smo}(M)$, the number of critical points of $f_{t}$ is constant with respect to $t$.
Proof.
Since $f_{t}$ is assumed to be Morse–Smale for any $t\in I$, each $f_{t}$ for any $t\in I$ has a finite number of isolated critical points and each $f_{t}$ induces a CW decomposition $\widetilde{M}$ with a $k$-cell for each critical point of index $k$ of $f_{t}$ of $M$, i.e. a CW complex which is homotopy equivalent to $M$. Thus, the Euler characteristic of the two agree, i.e. $\chi(M)=\chi(\widetilde{M})$. Since $\chi(\widetilde{M})$ is defined as the alternating sum over the number of $k$-cells of $\widetilde{M}$, this means that if any new critical points of $f_{t}$ appear with increasing $t$, they can only appear in pairs such that they leave the Euler characteristic unchanged.
Due to Lemma 4.5 (1) we can, without loss of generality, assume that there is a $t_{0}$ such that $f_{t_{0}}$ has two more non-degenerate critical points than $f_{t}$ for $t<t_{0}$. Let $c_{1},c_{2}$ be those two new critical points. We can chose disjoint neighborhoods $c_{1}\in U_{1}$, $c_{2}\in U_{2}$ and a small perturbation $\widetilde{f}_{t}$ such that, without loss of generality, $c_{1}$ appears at a time $t_{1}>t_{0}$ but $f_{t}$ and $\widetilde{f}_{t}$ agree outside of $U_{1}$.
Such a perturbation can be achieved as follows: Choose $\epsilon>0$ such that $B_{\epsilon}(c_{1})\subset U_{1}$ with respect to the metric on $M$ induced by the given Riemannian metric. We define the smooth function $\rho\colon[t_{0}-\delta,t_{0}+\delta]\rightarrow[t_{0}-\delta,t_{0}+\delta]$ by
$\rho(t)\coloneqq\begin{cases}\frac{t+t_{0}-\delta}{2}&\text{for }t_{0}-\delta\leq t<t_{0}\\ \frac{3t-t_{0}-\delta}{2}&\text{for }t_{0}\leq t\leq t_{0}+\delta\end{cases}.$
Then we define a perturbation
$\widetilde{f}(x,t)\coloneqq\begin{cases}f(x,t)&\text{for }x\in M\setminus B_{\epsilon}(c_{1}),\ t\in I\\ f(x,t)&\text{for }x\in B_{\epsilon}(c_{1}),\ t\in I\setminus[t_{0}-\delta,t_{0}+\delta]\\ f(x,(1-\frac{d(x,c_{1})}{\epsilon})\cdot\rho(t)+\frac{d(x,c_{1})}{\epsilon}\cdot t)&\text{for }x\in B_{\epsilon}(c_{1}),\ t\in[t_{0}-\delta,t_{0}+\delta]\end{cases},$
where the formulas are supposed to be interpreted for an appropriate choice of coordinates. Since $\widetilde{f}_{t}$ has no degenerate critical points by construction, $\widetilde{f}_{t}$ is a path of Morse functions and $\widetilde{f}_{t_{0}}$ can be arbitrarily closely approximated by a Morse–Smale function $f^{\prime}_{t_{0}}$.
But then, the function $f^{\prime}_{t_{0}}$ would induce a CW decomposition of $M$ with a different Euler characteristic which is a contradiction to the CW decomposition being homotopy equivalent to $M$. Thus, no new critical points can appear along $f_{t}$.
It can be proved analogously that along $f_{t}$ the number of critical points cannot decrease. All in all, the statement follows. ∎
Proof of Proposition 4.4.
It follows directly from Lemma 4.6 that all functions in $\mathcal{N}(f)$ have the same number of critical points. Since by Lemma 4.5 the eigenvalues of $Hf_{t}(p)$ depend continuously on $t$ for all $p\in M$, the indices of critical points cannot change along a path $f_{t}$ of Morse functions because otherwise, the intermediate value theorem implies that for some $t^{\prime}\in I$ there must have been a critical point with eigenvalue 0 which is a contradiction to $f_{t}$ being a path of Morse functions. Hence, the indices of the critical points remain unchanged along any path $f_{t}$. Thus, the statement follows. ∎
Corollary 4.7.
[…]
The result Corollary 4.7 can be extended to connected components of Morse functions if one does not fix a Riemannian metric, but allows the metric to be changed in order to make Morse functions Morse–Smale or works with choosing appropriate gradient-like flows instead. Both these approaches come with the problem that they do not give rise to a well-defined assignment of a CW complex to a Morse function because they involve non-canonical choices. This is the reason why we decided to fix a Riemannian metric at the cost of restricting to the space of Morse–Smale functions with respect to the chosen Riemannian metric. This way, we made the necessary non-canonical choice beforehand.
We continue with the investigation of the induced gradient-flows because they are responsible for the attaching maps of the CW decomposition induced by a Morse–Smale function.
To fix notation, for a Morse function $f\colon M\rightarrow\mathbb{R}$ on a Riemannian manifold $M$ with Riemannian metric $g$, we denote the gradient vector field of $f$ with respect to $g$ by $\nabla f$. Moreover, we denote by $\phi_{f}(x,\theta)$ the flow generated by $-\nabla f$. Let further $\operatorname{Cr}(f)$ denote the set of critical points of $f$.
Theorem 4.8.
[…]
Proof.
Let $f_{t}$ be a path from $f$ to $g$ in $\mathcal{MS}_{\text{smo}}(M)$. Since $t$ is a parameter with respect to differentiation, $\nabla f_{t}$ depends continuously on $t$. Since the flow $\phi_{t}(x,\theta)\colon M\times\mathbb{R}\rightarrow M$ of the gradient field $\nabla f_{t}$ is the solution to the initial value problem
$\begin{split}\frac{d}{d\theta}\phi_{t}(x,\theta)&=-(\nabla f_{t})(\phi_{t}(x,\theta))\\ \phi_{t}(x,0)&=x\end{split}$
and solutions to initial value problems depend smoothly on their initial values, the time-dependent flow $\phi_{t}(x,\theta)$ depends continuously on $t$. It is a standard trick in dynamics to replace a time-dependent flow $\phi_{t}$ by a time-independent one $\Phi$ as follows:
$\Phi\colon(M\times\mathbb{R})\times\mathbb{R}\rightarrow M\times\mathbb{R},\ \Phi((x,t),\theta)\coloneqq(\phi_{t}(x,\theta),t+\theta)$
It is a classical result that for any fixed $t$ and for any starting point $x\in M$ the limits $\lim\limits_{\theta\rightarrow\pm\infty}\phi_{t}(x,\theta)$ exist and are critical points of $f_{t}$. Since $\Phi$ is a continuous flow, this implies that the critical points which are connected by a given flow line also depend continuously on $t$. Together with Proposition 4.4, the statement follows. ∎
Corollary 4.9.
[…]
Corollary 4.10.
It follows from Theorem 4.8 that Cerf’s map from Proposition 4.2 extends to a map $\eta\colon\mathcal{M}(\mathcal{N})\rightarrow\mathcal{M}_{\mathcal{N}}$, where $\mathcal{M}_{\mathcal{N}}$ denotes the CW decomposition of $M$ induced by any Morse function in the connected component $\mathcal{N}$ of $M$. Moreover, Cerf’s proof of Proposition 4.2 shows that this instance of $\eta$ is a topological submersion and compatible with the respective stratifications, too.
Building on Corollary 4.9, we can extend the map $\eta$ from Proposition 4.2 to the connected component $\mathcal{N}(f)$ of a given Morse–Smale function $f$. By Proposition 4.2, it still follows that $\eta\colon\mathcal{N}(f)\rightarrow\mathcal{M}_{\mathcal{N}(f)}\coloneqq\mathcal{M}(M_{f})$ is a topological submersion which is compatible with the respective stratifications.
Using Cerf’s stratification of the space of Morse functions, especially the connection between different connected components given by cancellations and creations of pairs of critical points, we want to further extend $\eta$ to certain systems of connected components:
Definition 4.11.
Consider the set $\pi_{0}(\mathcal{MS}_{smo}(M))$ of connected components of the space of Morse–Smale functions on $M$. Let $\mathcal{N}_{1},\mathcal{N}_{2}\in\pi_{0}(\mathcal{MS}_{smo}(M))$ be connected components. Then we define a partial order $\leq$ on $\pi_{0}(\mathcal{MS}_{smo}(M))$ generated by $\mathcal{N}_{1}<\mathcal{N}_{2}$ if for any $f\in\mathcal{N}_{2}$ there is a $g\in\mathcal{N}_{1}$ such that $g$ arises from $f$ by a cancellation of a pair of critical points. We call $\leq$ the cancellation order on $\pi_{0}(\mathcal{MS}_{smo}(M))$.
Remark 4.12.
It is shown in [Cer70, Section 3] that if two Morse functions are related to each other by a cancellation of one pair of critical points, then all paths between them are homotopic to one which traverses the first stratum $\mathcal{F}^{1}$ in one isolated point. Hence, we can think about two connected components $\mathcal{N}_{1},\mathcal{N}_{2}$ such that $\mathcal{N}_{1}<\mathcal{N}_{2}$ as being neighbors, only separated by a codimension one stratum which corresponds to a hyperplane in $\mathcal{M}(\mathcal{N}_{2})$ via $\eta$.
Proposition 4.13.
@@
Let $f,g\in\mathcal{MS}_{smo}(M)$ be two Morse–Smale functions such that $f$ arises from $g$ by a cancellation of one pair of critical points. Then $M_{f}$ and $M_{g}$ are simply homotopy equivalent by a simple collapse $M_{g}\rightarrow M_{f}$/ a simple extension $M_{f}\rightarrow M_{g}$ of the two cells which correspond to the canceled critical points at the corresponding sublevel complex. In particular, the map $M_{f}\rightarrow M_{g}$ is non-degenerate.
@@
Let $\mathcal{N}_{1}<\mathcal{N}_{2}\in\pi_{0}(\mathcal{MS}_{smo}(M))$ be connected components which are related by cancellations of critical points. Then $\mathcal{A}(M_{\mathcal{N}_{1}})$ canonically embeds into $\mathcal{A}(M_{\mathcal{N}_{2}})$ by Lemma 3.23. Moreover, we have an induced map $\mathcal{M}(M_{f})\rightarrow\mathcal{M}(M_{g})$ given by Definition 3.26.
5. Relationship to other concepts in persistent topology
@@ 5.1. The space of discrete Morse matchings
In the literature, there is already work on a so-called space, or complex, of discrete Morse functions. In fact, this name has been used with a slight abuse of terminology since these works refer to a complex whose simplices correspond to Morse matchings rather than Morse functions.
@@
Let $X$ be a simplicial complex. The complex of discrete Morse matchings $\mathfrak{M}(X)$ has as vertices cover relations in $D(X)$ and simplices sets of cover relations in $D(X)$ that correspond to acyclic matchings on $D(X)$.
@@
Let $X$ be a CW complex. The complex of discrete Morse matchings $\mathfrak{M}(X)$ on $X$ has vertices cover relations in $D(X)$ that correspond to regular faces and simplices sets of vertices that correspond to acyclic matchings on $D(X)$.
Since the geometric structure of the simplices, i.e. their points, do not have any obvious meaning for discrete Morse matchings, we will not distinguish between $\mathfrak{M}(X)$ and its face poset $D(\mathfrak{M}(X))$ in our notation.
@@
The minimal elements of both posets are in canonical bijection because in both cases they correspond to the cover relations in $D(X)$ that correspond to regular face relations. Then collections of cover relations that form an acyclic matching in $\mathfrak{M}(X)$ are mapped to intersections of hyperplanes in $\mathcal{L}(\mathcal{A})$ that correspond to the same acyclic matching. ∎
@@ 5.2. The space of merge trees
In order to give a model for the moduli space of merge trees, we introduce a notion of edit morphisms to relate different combinatorial merge trees to each other.
@@ 1. The identity:
    The poset $P$ is left unchanged.
@@ 2. Adding a leaf:
    A new minimal element $x$ is added in one of two ways: either, we just introduce the relation $x\leq y$ for a pre-existing non-minimal element $y\in P$, or we add a new non-minimal element $y^{\prime}$ between two pre-existing elements $z_{1}\leq y\leq z_{2}$ and add the relation $x\leq y$.
@@ 3. Removing a leaf:
    A pre-existing minimal element $x\in P$ is removed. If $x$s parent node $y$ then only has one child node $x^{\prime}$, we also remove $y$.
  4. Splitting an inner node
    If a node $z\in P$ has at least 3 child nodes, $z$ can be split into two nodes $z_{1}\leq z_{2}$ such that there is no $x\in P$ with $z_{1}\leq x\leq z_{2}$. The previous child nodes of $z$ are then distributed among $z_{1}$ and $z_{2}$ such that $z_{1}$ gets at least two child nodes and $z_{2}$ gets at least one child node other than $z_{2}$.
@@ 5. Merging two adjacent inner nodes:
    If there are two non-minimal elements $z_{1}\leq z_{2}\in P$ such that there is no element between them, i.e. there exists no $x:z_{1}\leq x\leq Z_{2}$, we construct the quotient $P/_{z_{1}=z_{2}}$. This creates an element $z$ instead of $z_{1}$ and $z_{2}$ such that all previous child nodes of $z_{1}$ and $z_{2}$ become child nodes of $z$.
An edit morphism $\eta\colon P\rightarrow P^{\prime}$ between posets of merge tree type is a composition of edit moves together with an automorphism of $P^{\prime}$. An edit morphism is called elementary if it performs only one of the edit moves. An edit morphism is called trivial if it has the identity as its only corresponding edit move.
@@
Let $P,P^{\prime}$ be posets of merge tree type such that $P^{\prime}$ arises from $P$ by either an addition of a leaf or a splitting of an inner node. Then any corresponding edit morphism induces a poset morphism $\eta\colon P\rightarrow P^{\prime}$ given by the inclusion of the preexisting nodes.
We identify spaces of discrete functions, as introduced in Definition 3.1, with sets of set maps $\operatorname{Map}(P,\mathbb{R})$ for posets $P$ of CW type. Moreover, we identify sets of merge trees with the same underlying combinatorial merge tree $P$ with sets of order-preserving maps $\operatorname{Map}_{\leq}(P,\mathbb{R})$ for posets $P$ of merge tree type.
@@
$Mer\coloneqq\coprod\limits_{P\text{ of merge tree type}}\operatorname{Map}_{\leq}(P,\mathbb{R}).$
For a merge tree $\theta\in Mer$ we denote the underlying combinatorial merge tree, i.e. the source of $\theta$, by $s(\theta)$.
@@
$Mer_{st}\coloneqq\coprod\limits_{P\text{ of merge tree type}}\operatorname{Map}_{<}(P,\mathbb{R}),$
@@
$Mer_{wb}\coloneqq\{\theta\in Mer_{st}|\forall y\in s(\theta)\text{ the restriction }\theta_{\lvert s(\theta)_{\leq y}}\text{ has a minimum.}\},$. } ,
@@
In order to ensure that $Mer$ is actually a set, we would need to fix a way of how to turn the category of finite sets into a small category. Once that is done, we can do the same for posets since there are only finitely many relations, in particular finitely many partial orders, on any finite set. Then posets of merge tree type form a subset of said set and the disjoint union above is actually indexed over a set.
@@
Let $X$ be a regular CW complex. We define a map $M\colon\mathcal{M}(X)\rightarrow Mer$ as follows: Let $f\in\mathcal{M}(X)$ be a discrete Morse function and let $\operatorname{Cr}_{i}(f)$ the set of critical cells of dimension $i$ of $f$. We create a poset of merge tree type $P$ as follows: for each critical vertex $\sigma\in\operatorname{Cr}_{0}(f)$ we add an element $\sigma$ to $P$. For each time connected components $\sigma_{1},\dots,\sigma_{k}$ of sublevel complexes merge within the filtration, we add an element $\tau$ to $P$ and add the relations $\sigma_{i}\leq\tau$. The elements $\tau$ correspond to certain critical edges $\tau\in\operatorname{Cr}_{1}(f)$. Then we define $M(f)(\sigma)\coloneqq f_{\sigma}$, i.e. the elements of $P$ get their functions values under $M(f)$ from the critical values of the corresponding critical cells.
@@
We proceed by considering how crossing hyperplanes of $\mathcal{A}$ and $\mathcal{H}$ relates to edit moves between the induced combinatorial merge trees.
@@
It is well known, e.g. [For98, Theorem 11.1], that whenever a Morse function $f\in\mathcal{M}(X)$ induces a unique gradient path between a critical $d$ cell $\sigma$ and a critical $d+1$ cell $\tau$ such that all face relations along the gradient path are regular, the gradient path can be inverted, matching $\sigma$ and $\tau$ in the process. If $\sigma$ and $\tau$ have consecutive function values, then $\sigma$ is a face of $\tau$ and we are in the situation of Lemma 5.10. Otherwise, there are other cells with function values between $f(\sigma)$ and $f(\tau)$ and we have to cross hyperplanes of $\mathcal{A}$ corresponding to all face relations along said unique gradient path between $\sigma$ and $\tau$ as well as hyperplanes of $\mathcal{H}\setminus\mathcal{A}$ corresponding to cells that are not contained in the gradient path with function values between $f(\sigma)$ and $f(\tau)$. It should be remarked that during this procedure, we cannot cross all hyperplanes of $\mathcal{A}$ corresponding to the face relations along the unique gradient path simultaneously because otherwise we would leave the space of discrete Morse functions.

### v2

@@ 4.1. Smooth Morse Theory and Cerf Theory
Let $f,g\in\mathcal{MS}(M)$ be two Morse–Smale functions such that $f$ arises from $g$ by a cancellation of one pair of critical points. Then $M_{f}$ and $M_{g}$ are simply homotopy equivalent by a simple collapse $M_{g}\rightarrow M_{f}$/ a simple extension $M_{f}\rightarrow M_{g}$ of the two cells which correspond to the canceled critical points at the corresponding sublevel complex.
@@
Let $\mathcal{N}_{1}<\mathcal{N}_{2}\in\pi_{0}(\mathcal{MS}(M))$ be path components, which are related by cancellations of critical points. Then $\mathcal{A}(M_{\mathcal{N}_{1}})$ canonically embeds into $\mathcal{A}(M_{\mathcal{N}_{2}})$ and we have an embedding $\mathcal{M}(M_{\mathcal{N}_{1}})\rightarrow\mathcal{M}(M_{\mathcal{N}_{2}})$ by Proposition 3.41.
Remark 4.14.
Although the map $\eta$ from Corollary 4.9 is only defined on the path components of the space of Morse–Smale functions, Corollary 4.13 allows us to relate the spaces of discrete Morse functions on CW decompositions induced from different path components to each other. Provided one finds a good notion of space of discrete Morse functions on a simple homotopy type, as we have conjectured in Remark 3.42, one could “globalize” the map $\eta$ and extend it to a topological submersion $\eta\colon\mathcal{MS}(M)\rightarrow\mathcal{M}(\mathcal{SH}(M))$, where $\mathcal{SH}(M)$ denotes the simple homotopy type of all CW decompositions of $M$.
4.1.1. Discrete Cerf theory
Since one of the central works that inspired this one is Cerf’s structural description of the geometric interpretation of cancellation moves in the space of Morse functions, see [CER70], we want to provide a brief translation of this theory to the much easier discrete case.
Smooth case Discrete case
-stratification by codimension -stratification by braid arrangement
-function has a critical point of codim 1 -function lies inside a hyperplane of the Morse arrangement
-function has a critical point of higher codimension -function lies inside the intersection of hyperplanes whose cover relations share a cell
-universal unfolding of a critical point -basis of the orthogonal complement of the (intersection of) hyperplane(s)
-cancellation of a pair of critical points -cancellation of a pair of critical cells
-creation of a pair of critical points -removing a matched pair from the Morse matching (potential need for an internal extension)
-function has a critical value of higher codimension -function lies in the intersection of hyperplanes that belong to the braid arrangement but not to the Morse arrangement
-moving critical points past each other -crossing a hyperplane that belongs to the braid arrangement but not to the Morse arrangement
Figure 5. Comparison between smooth and discrete Cerf theory
The well-known Cerf diagrams apply the same way in the discrete case as they do in the smooth case. Moreover, the Cerf diagrams can in the discrete case quite explicitly illustrate the corresponding path of discrete functions, seen inside the correct orthogonal projection of the space of discrete functions.
The central difference between the two theories is, that the creation of pairs of critical points is always available in the smooth case, whereas in the discrete case the CW complex at hand may be lacking the necessary cells. In order to fix this, one can always perform an internal extension, but this technically changes the space of discrete functions under consideration. In order to truly fix this, one would need to make sense of the notion of space of discrete Morse functions on a simple homotopy type, as we conjectured in Remark 3.42.
@@ 4.2. The space of discrete Morse matchings
In the literature, there is already work on a notion of complex of discrete Morse functions. In fact, this name has been used with a slight abuse of terminology since these works refer to a complex whose simplices correspond to Morse matchings rather than Morse functions. In order to make this difference clear, we use the different name “complex of discrete Morse matchings” to refer to this concept.
@@
Let $X$ be a simplicial complex. The complex of discrete Morse matchings $\mathfrak{M}(X)$ has as vertices cover relations in $F(X)$ and simplices sets of cover relations in $F(X)$ that correspond to acyclic matchings on $D(X)$.
@@
Let $X$ be a CW complex. The complex of discrete Morse matchings $\mathfrak{M}(X)$ on $X$ has vertices cover relations in $F(X)$ that correspond to regular faces and simplices sets of vertices that correspond to acyclic matchings on $D(X)$.
Since the geometric structure of the simplices, i.e. their points, do not have any obvious meaning for discrete Morse matchings, we will not distinguish between $\mathfrak{M}(X)$ and its face poset $F(\mathfrak{M}(X))$ in our notation.
@@
The minimal elements of both posets are in canonical bijection because in both cases they are defined by the regular cover relations in $F(X)$. Then collections of cover relations that form an acyclic matching in $\mathfrak{M}(X)$ are mapped to intersections of hyperplanes in $\mathcal{L}(\mathcal{A})$ that correspond to the same acyclic matching. ∎
@@ 4.3. The space of merge trees
In order to give a model for the space of merge trees, we introduce the notions of edit moves and edit morphisms to relate different combinatorial merge trees to each other.
@@ 1. The identity:
    The poset $P$ is left unchanged. We denote this edit move by $\operatorname{id}$.
@@ 2. Adding a leaf:
    A new minimal element $x$ is added in one of two ways: either, we just introduce the relation $x\leq y$ for a pre-existing non-minimal element $y\in P$, or we add a new non-minimal element $y^{\prime}$ between two pre-existing elements $z_{1}\leq y^{\prime}\leq z_{2}$ and add the relation $x\leq y^{\prime}$. As an exceptional case, if $P$ is the one element poset $\{r\}$, in particular, there are no inner nodes, we allow to add two leaves $x<r,y<r$ simultaneously. We denote this edit move by $\operatorname{Add}(x,y)$.
@@ 3. Removing a leaf:
    A pre-existing minimal element $x\in P$ is removed. If $x$’s parent node $y$ has only the one child node $x$, we remove $y$ in addition to $x$. We denote this by $\operatorname{Rem}(x,y)$.
  4. Splitting an inner node  [^16] On combinatorial merge trees, this type of edit move provides multiple different options of edit moves depending on how the child nodes are precisely being distributed among $z_{1}$ and $z_{2}$.
    If a node $z\in P$ has at least 3 child nodes, then $z$ can be split into two nodes $z_{1}\leq z_{2}$ such that there is no $x\in P$ with $z_{1}\leq x\leq z_{2}$. The previous child nodes of $z$ are then distributed among $z_{1}$ and $z_{2}$ such that $z_{1}$ gets at least two child nodes and $z_{2}$ gets at least one child node other than $z_{1}$. We denote this edit move by $\operatorname{Spl}(z,z_{1},z_{2})$.
@@ 5. Merging two adjacent inner nodes:
    If there are two non-minimal elements $z_{1}\leq z_{2}\in P$ such that $z_{2}$ covers $z_{1}$, i.e. $z_{1}<z_{2}$ and there exists no $x:z_{1}\leq x\leq z_{2}$, we construct the quotient $P/_{z_{1}=z_{2}}$. This creates an element $z$ instead of $z_{1}$ and $z_{2}$ such that all previous child nodes of $z_{1}$ and $z_{2}$ become child nodes of $z$. We denote this edit move by $\operatorname{Mer}(z_{1},z_{2},z)$.
An edit morphism $\eta\colon P\rightarrow P^{\prime}$ between posets of merge tree type is a composition of edit moves composed with an automorphism of $P^{\prime}$. An edit morphism is called elementary if it performs only one of the edit moves. An edit morphism is called trivial if it has the identity as its only corresponding edit move. We some times abbreviate the notion elementary edit morphism by EEM. By slight abuse of notation, we denote elementary edit morphisms by their corresponding edit move.
@@
Let $P,P^{\prime}$ be posets of merge tree type such that $P^{\prime}$ arises from $P$ by either an addition of a leaf or a splitting of an inner node. Then any corresponding edit morphism induces a poset morphism $\eta\colon P\rightarrow P^{\prime}$ given by the inclusion of the preexisting nodes and mapping nodes $x$ that are being split to maximal one $z_{2}$ of the new nodes.
Additionally, we want to remark that for each finite poset of merge tree type $P$, there is only a finite number of elementary edit moves that can be applied to $P$. Those are:
  1. (1)
    Adding one leaf to any inner node or edge,
  2. (2)
    Removing at most one leaf per leaf (in general not possible for all leaves.)
We identify spaces of discrete functions, as introduced in Definition 3.1, with sets of set maps $\operatorname{Map}(P,\mathbb{R})$ for posets $P$ of CW type. Moreover, we identify sets of merge trees with the same underlying combinatorial merge tree $P$ with sets of order-preserving maps $\operatorname{Map}_{\leq}(P,\mathbb{R})$ for posets $P$ of merge tree type. We use the following model [^17] By model we mean a set that has one specific poset of merge tree typ for every isomorphism class of posets of merge tree type. for the class of posets of merge tree type:
Definition 4.20.
We construct the set of combinatorial merge trees $Mer_{comb}$ as follows: Let $n\in\mathbb{N}_{\geq 1}$. We first define a set of representatives for posets of merge tree type with $n$ nodes, which we denote by $Mer_{comb}^{n}$: we consider the set of nodes $\{1,2,\dots,n\}$ and fix a collection of poset structures on $\{1,2,\dots,n\}$ such that there is one representative for each isomorphism type of posets of merge tree type with $n$ nodes. This is possible because the collection of partial orders on $\{1,\dots,n\}$ is a subset of the set of relations $\{1,\dots,n\}\times\{1,\dots,n\}$, so in particular there is only a finite amount of those. We define
$Mer_{comb}\coloneqq\coprod\limits_{n\in\mathbb{N}_{\geq 1}}Mer_{comb}^{n}.$
We do not claim that the choice of representatives in $Mer_{comb}$ is canonical in any way - it is not. However, this is fine for our purposes. From now on, we will not address this choice of representatives any further: we will always assume that composition with an additional poset isomorphism from the target of any morphism under consideration to one of the chosen representatives takes care of all such issues.
@@
$Mer\coloneqq\coprod\limits_{P\in Mer_{comb}}\operatorname{Map}_{\leq}(P,\mathbb{R}).$
For any merge tree $\theta\in Mer$ we denote the underlying combinatorial merge tree, i.e. the source of $\theta$, by $s(\theta)$.
@@
$Mer_{st}\coloneqq\coprod\limits_{PMer_{comb}}\operatorname{Map}_{<}(P,\mathbb{R}),$
@@
$Mer_{wb}\coloneqq\{\theta\in Mer_{st}|\ \forall y\in s(\theta)\text{ the restriction }\theta_{\lvert s(\theta)_{\leq y}}\text{ has a minimum}\},$
@@
In the definition above we slightly abuse notation by identifying a merge tree with its height function. Since any height function $\theta$ implicitly also contains the information of the underlying combinatorial merge tree as the domain of $\theta$, this should hopefully not lead to any confusion.
We consider how edit morphisms can be used to shift height functions from one combinatorial merge tree to another:
@@
We say that a collection of edit morphisms $\{\eta_{i}\}_{i\in I}$ is/goes in the same direction if there is a way to compose all of the $\eta_{i}$ to one morphism and all $\{\eta_{i}\}_{i\in I}$ are either additions of leaves and splittings of inner nodes, or removals of leaves and merges of inner nodes.
Let $\eta$ be a composition of edit morphisms in the same direction that consists of additions of leaves and splittings of inner nodes. We denote by $\operatorname{L}(\eta)$ the collection of pairs $(x,y)$ such that $x$ is a leave that is being added by $\eta$ to its parent node $y$. Moreover, we denote by $\operatorname{S}(\eta)$ the collection of pairs of nodes $(x,x_{i})$ such that $x_{i}$ was created by $\eta$ splitting $x_{i}$ off of $x$.
In the case of merge trees whose underlying combinatorial merge trees are related by compositions of elementary edit morphisms in the same direction, there is a convenient way of shifting height function values between combinatorial merge trees:
Definition 4.24.
Let $\theta\in\operatorname{Map}_{\leq}(P,\mathbb{R})$ be a merge tree $P$ and let $P^{\prime}$ be a poset of merge tree type such there is a composition of EEMs in the same direction $\eta\colon P\rightarrow P^{\prime}$.
Then by Remark 4.19 we have an induced injective morphism $\eta\colon P\rightarrow P^{\prime}$ if $\eta$ is a composition of adding leaves and splitting inner nodes. We extend the inclusion $\eta\colon P\rightarrow P^{\prime}$ associated to any composition of adding leaves and splitting inner nodes to a map $\eta_{*}\colon\operatorname{Map}_{\leq}(P,\mathbb{R})\times\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})\rightarrow\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$ by
[…]
Remark 4.25.
If $\eta$ is a composition of removing leaves and merging inner nodes, we have an induced inclusion $\eta\colon P^{\prime}\rightarrow P$ in the other direction by the same construction for the inverse composition of EEMs in the same direction.
Definition 4.26.
Let $X$ be a regular CW complex. We define a map $M\colon\mathcal{M}(X)\rightarrow Mer$ as follows: Let $f\in\mathcal{M}(X)$ be a discrete Morse function and let $\operatorname{Cr}_{i}(f)$ the set of critical cells of dimension $i$ of $f$. We create a poset of merge tree type $\tilde{P}$ as follows: for each critical vertex $\sigma\in\operatorname{Cr}_{0}(f)$ we add an element $\sigma$ to $\tilde{P}$. For each time when connected components $\sigma_{1},\dots,\sigma_{k}$ of sublevel complexes merge within the filtration, we add an element $\tau$ to $\tilde{P}$ and add the relations $\sigma_{i}\leq\tau$. The elements $\tau$ correspond to certain critical edges $\tau\in\operatorname{Cr}_{1}(f)$. Moreover, we identify $\tilde{P}$ with the unique representative $P\in Mer_{comb}$ of the isomorphism type of $\tilde{P}$. Then we define $M(f)(\sigma)\coloneqq f_{\sigma}$, i.e. the elements of $P$ get their function values under $M(f)$ from the critical values of the corresponding critical cells.
Example 4.27.
We consider an example of a discrete Morse function on a regular CW complex and its induced merge tree:
$\bullet\bullet$0223415450134230110
The colors indicate which critical cells correspond to which nodes of the merge tree. Here we have drawn the merge tree as its Hasse diagram.
@@
We proceed by considering how crossing hyperplanes of $\mathcal{A}$ and of $\mathcal{H}$ relates to edit moves between the induced combinatorial merge trees.
@@
It is well known, e.g. [FOR98, Theorem 11.1], that whenever a Morse function $f\in\mathcal{M}(X)$ induces a unique gradient path between a critical $d$ cell $\sigma$ and a critical $d+1$ cell $\tau$ such that all face relations along the gradient path are regular, the gradient path can be inverted, matching $\sigma$ and $\tau$ as a result. If $\sigma$ and $\tau$ have consecutive function values, then $\sigma$ is a face of $\tau$ and we are in the situation of Lemma 4.29. Otherwise, there are other cells with function values between $f(\sigma)$ and $f(\tau)$ and we have to cross hyperplanes of $\mathcal{A}$ corresponding to all face relations along said unique gradient path between $\sigma$ and $\tau$ as well as hyperplanes of $\mathcal{H}\setminus\mathcal{A}$ corresponding to cells that are not contained in the gradient path with function values between $f(\sigma)$ and $f(\tau)$. It should be remarked that during this procedure, we cannot cross all hyperplanes of $\mathcal{A}$ corresponding to the face relations along the unique gradient path simultaneously, if the path is of length larger than 1, because otherwise we would leave the space of discrete Morse functions.